OpenStudy (juscallmesteve):

Differentiate the problem using Logarithmic Differentiation \[y=(x+6)^x\]

OpenStudy (juscallmesteve):

When I do this I get \[y \prime=(x+6)^x(\frac{ x }{ (x+6) }+\ln(x+6))\]

OpenStudy (juscallmesteve):

also how do i give people medals for helping me?

OpenStudy (solomonzelman):

To give user a medal, click Best Response next to their reply. Now, as far as the differentiation process here goes...

OpenStudy (juscallmesteve):

O i did not know that game someone a medal

OpenStudy (solomonzelman):

In general, \(\color{#000000}{ \displaystyle y=f(x)^{g(x)} }\) \(\color{#000000}{ \displaystyle \ln y=\ln \left(f(x)^{g(x)}\right) }\) \(\color{#000000}{ \displaystyle \ln y=g(x)\ln \left(f(x)\right) }\) differentiate both sides \(\color{#000000}{ \displaystyle \frac{y'}{y}=g'(x)\ln \left(f(x)\right)+g(x)\frac{f'(x)}{f(x)} }\) \(\color{#000000}{ \displaystyle y'=y\cdot \left(g'(x)\ln \left(f(x)\right)+g(x)\frac{f'(x)}{f(x)} \right) }\) \(\color{#000000}{ \displaystyle y'=f(x)^{g(x)} \cdot \left(g'(x)\ln \left(f(x)\right)+g(x)\frac{f'(x)}{f(x)} \right) }\) this is the process in general.

OpenStudy (solomonzelman):

and your result is correct.

OpenStudy (solomonzelman):

(so I guess there wasn't a need going over that process once again)

OpenStudy (juscallmesteve):

Thank you very much for all of that information

OpenStudy (solomonzelman):

Not a problem!

OpenStudy (solomonzelman):

good job!